OpenStudy (anonymous):

1. Choose the expression that forms an equation with the given expression. 24 – 5 X 2 = (5 + 3) • 2 2(6 + 1) 5 • 8 – 2 4 • 9 + 2 Question 2.2. Choose the expression that completes the equation. 30 ÷ 2 + 10 = ________ 3(2 + 3) 7(2 + 1) 6(3 + 7) 5(4 +1) Question 3.3. Determine which value makes the open sentence true. 12 + 3x = 30 x = 2 x = 4 x = 5 x = 6

OpenStudy (anonymous):

Ill help! I love these.

OpenStudy (anonymous):

3 is 6 but I cnat tell what the others say

OpenStudy (anonymous):

Ok,Thank you so much

OpenStudy (anonymous):

1. Choose the expression that forms an equation with the given expression. 24 – 5 x 2 = (5 + 3) x 2 2(6 + 1) 5 x 8 – 2 4 x 9 + 2 Question 2.2. Choose the expression that completes the equation. 30 ÷ 2 + 10 = ________ 3(2 + 3) 7(2 + 1) 6(3 + 7) 5(4 +1)

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

Can you help me with the one above

OpenStudy (anonymous):

Ugh...Those triangles are making it hard to understand lol

OpenStudy (anonymous):

Can you tell me it in word form? Lol

OpenStudy (anonymous):

1. Choose the expression that forms an equation with the given expression. 24 - 5 times 2 = (5 + 3) times 2 2(6 + 1) 5 times 8 – 2 4 times 9 + 2 Question 2.2. Choose the expression that completes the equation. 30 divided 2 + 10 = 3(2 + 3) 7(2 + 1) 6(3 + 7) 5(4 +1)

OpenStudy (anonymous):

Thats better!

OpenStudy (anonymous):

B and D

OpenStudy (anonymous):

@tmcfaddin

OpenStudy (anonymous):

5. Determine which value makes the open sentence true. 1.2 = 3x (Points : 1) x = 0.4 x = 3.6 x = 2.5 x = 1.8

OpenStudy (anonymous):

That's the last one I need help with